Good day,

2COF2(g)⇌CO2(g)+CF4(g), Kc=4.30

If only COF2 is present initially at a concentration of 2.00 M, what concentration of COF2
remains at equilibrium?
Express your answer with the appropriate units.
Thanks
1
Expert's answer
2020-08-09T08:42:57-0400

Kc=[CO2][CF4]/[COF2]2

let [CO2] = [CF4] = x

then [COF2] = 2.00 - 2x

"\frac{x^{2}}{(2.00-2x)^{2}} = 4.30"

x=0.806

concentration of [COF2] = 2.00 - 2x = 0.388 M